# AI Competitions and Benchmarks: towards impactful challenges with post-challenge papers, benchmarks and other dissemination actions

### v2

@@ 4.4.5 Deeper analysis of the submissions
For example, Fig. 4.8 shows the performance of different submissions for different tasks.
Figure 4.8: When a competition involves different tasks or different underlying metrics, a radar plot can provide a nice visual comparison of the submissions along all those dimensions as CrosbyBSHCH19 which tested agents on several tasks.
Another example, in Fig.4.9 shows the performance of some of the best participants in a clustering competition (same participants as in Fig.4.7). Although the participants had to optimise a single score, domain experts were satisfied to see these graphs showing that the best submissions maximise their cluster-finding algorithm’s robustness (with respect to ground truth parameters). One notable exception is #$\#100$, the only one showing a rising contribution in the bottom left graph. It indicated it was accidentally optimised for abnormal clusters, yielding a poor overall score, which was still interesting to domain experts.
Figure 4.9: Cluster finding probability as a function of six ground truth cluster 3D shape and dimension parameters, from the TrackML Accuracy competition TrackMLAccuracy2019
Pipeline. A challenge solution usually consists of a pipeline of steps, e.g., data preprocessing, feature engineering, model training, hyperparameter selection, ensemble, etc. It is super interesting to investigate step by step the choices available from participants and ablation study the contribution of options. Such a study is not trivial because, firstly, we need to split the steps of solutions, which is not necessarily logically clear; secondly, we need to modify many solutions to evaluate the options. Once an ablation study is done, we could further ensemble the best options and output a revised version.
Ensembling. For each submission, the participant made a decision for each step of the pipeline. For some challenge types (i.e. classification), it should be possible to automatically ensemble different submissions to see if something could be gained from combining them Kgl2018TheRF . In general, human ensembling (mix and match of the various decisions taken at each step) can bring insights and improvements over even the very best submission. For example, one participant has built clever features but could have used them better.
@@
4.4.6 Description of the most interesting submissions
If one paper is written, which hosts contributions from the most interesting submissions (not necessarily the absolute best but also the ones deemed most originals), this section would hold sub-sections, each describing a submission. It is best that participants write themselves these sub-sections, following guidelines or a template provided by the organisers. Otherwise, the organisers can write themselves one or more subsections based on the material available to them.
If the participants have written separate papers, only brief summaries would be necessary here.
4.4.7 Scientific outcome
@@
4.4.8 Lessons on the challenge organisation itself
@@ 4.4.9 Conclusion
4.5 Post challenge benchmark
Benchmarks differ from competitions in many ways, as summarised in Table 4.1. We organise a competition to crowdsource a task and harvest the winning solutions. This competition usually lasts a couple of months and has multiple phases (public, private, etc.). We intentionally rank participants linearly due to their competitive nature, and people are not allowed to share the code directly. However, for benchmarks, we are interested in a research task and would like to invite people worldwide to contribute continuously. The benchmarks last much longer than competitions, usually never-ending, and only one phase is associated. Another big difference is that benchmarks encourage people to share ideas, solutions, code, and findings as much as possible because the goal is to push forward this research task. Thus, rich publications, seminars, and workshops are expected for communication.
By turning a challenge into a benchmark, we gain multiple benefits for different people. For challenge organisers, they give more time (possibly never-ending) to people around the world to continuously join the benchmark and make submissions to push forward the research task. For participants, they have more time and possibly can open-source datasets for their own research, and the leaderboard of research gives credit to their methods. For the platform of benchmarks, it is always better to have high-quality benchmarks and attract more people.
@@ 4.5 Post challenge benchmark
Table 4.1: Comparisons of competition and benchmark codabench .
Competition
Benchmark
Purpose
Crowdsourcing problems in a short time and harvesting solutions
Continuous fair evaluation over a long time period, in a unified framework
Phases
Multiple phases
Single-phase
Time period
Usually limited
Often never-ending
Cooperation & information sharing
Limited due to the competitive nature
As extensive as possible
Submissions
Usually algorithm predictions or algorithm code
Algorithm code or datasets; code or dataset name, description, documentation meta-data and/or fact-sheets; scoring programs for custom analyses
Outcome
leaderboard with usually a single global ranking based on one score from each team (last or best)
Table with all the submissions made; sorting with multiple scores possible; multiple analyses, graphs, figures, code sharing
4.6 Conclusion
In the chapter, we have covered the main actions in favour of the long-lasting impact of a challenge, in particular, with a post-challenge paper template and advice on how to turn the challenge into a benchmark.
@@ 4.6 Conclusion
inline,color=brown!40]I commend the authors on a good job at sketching the need for specific activities (such as a paper and benchmarks) for the competition to maximise its impact. Overall, the chapter is in good shape, but it can benefit from some extra work.
@@
Finally, in my personal view, the chapter reads more like a recipe of how to (best represented by the itemised lists of ingredients) than like a piece of work in which the authors share the accumulated wisdom in the field of how to maximise impact with post-challenge activities. In other words, I was hoping for more depth to accompany the breadth. I feel it would be useful, though I leave it to the discretion of the authors, to discuss examples of what different competitions do similarly and differently in terms of post-challenge activities, and how those differences shed light on what works and what doesn’t. In the paper template section, it would be useful to cite references with examples where papers report things in slightly divergent ways (which is good), and papers with sections that are particularly outstanding (e.g., scientific outcomes, ensembling, etc.) in the post-challenge papers in different domains. I know that this may be a tall order, but I honestly believe it will improve the chapter considerably. – Gustavo

### v5

@@ 4.5 Deeper analysis of the submissions
For example, Fig. 8 shows the performance of different submissions for different tasks.
Figure 8: When a competition involves different tasks or different underlying metrics, a radar plot can provide a nice visual comparison of the submissions along all those dimensions as (Crosby et al., 2019) which tested agents on several tasks.
Another example, in Fig. 9 shows the performance of some of the best participants in a clustering competition (same participants as in Fig. 7). Although the participants had to optimise a single score, domain experts were satisfied to see these graphs showing that the best submissions maximise their cluster-finding algorithm’s robustness (concerning ground truth parameters). One notable exception is #100#100\#100# 100, the only one showing a rising contribution in the bottom left graph. It indicated it was accidentally optimised for abnormal clusters, yielding a poor overall score, which was still interesting to domain experts.
Figure 9: Cluster finding probability as a function of six ground truth cluster 3D shape and dimension parameters, from the TrackML Accuracy competition (Amrouche et al., 2019)
Pipeline. A challenge solution usually consists of a pipeline of steps, e.g., data preprocessing, feature engineering, model training, hyperparameter selection, ensemble, etc. It is super interesting to investigate step by step the choices available from participants and ablation study the contribution of options. Such a study is not trivial because we need to split the steps of solutions, which is not necessarily logically clear; secondly, we need to modify many solutions to evaluate the options. Once an ablation study is done, we could further ensemble the best options and output a revised version.
Ensembling. For each submission, the participant made a decision for each pipeline step. For some challenge types (i.e. classification), it should be possible to automatically ensemble different submissions to see if something could be gained from combining them (Kégl et al., 2018). In general, human ensembling (mixing and matching the various decisions taken at each step) can bring insights and improvements to even the very best submission. For example, one participant has built clever features but could have used them better.
@@
4.6 Description of the most interesting submissions
If one paper is written, which hosts contributions from the most interesting submissions (not necessarily the absolute best but also the ones deemed original), this section would hold sub-sections, each describing a submission. Participants should write these sub-sections themselves, following guidelines or a template provided by the organisers. Otherwise, the organisers can write themselves one or more subsections based on the material available to them.
If the participants have written separate papers, only summaries would be necessary here.
4.7 Scientific outcome
@@
4.8 Lessons on the challenge organisation itself
@@ 4.9 Conclusion
5 Post challenge benchmark
Benchmarks differ from competitions in many ways, as summarised in Table 1. We organise a competition to crowdsource a task and harvest the winning solutions. This competition usually lasts a couple of months and has multiple phases (public, private, etc.). We intentionally rank participants linearly due to their competitive nature, and people are not allowed to share the code directly. However, for benchmarks, we are interested in a research task and would like to invite people worldwide to contribute continuously. The benchmarks last much longer than competitions, usually never-ending, and only one phase is associated. Another big difference is that benchmarks encourage people to share ideas, solutions, code, and findings as much as possible because the goal is to push forward this research task. Thus, rich publications, seminars, and workshops are expected for communication.
By turning a challenge into a benchmark, we gain multiple benefits for different people. Challenge organisers give people around the world more time (possibly never-ending) to join the benchmark and make submissions to push forward the research task. Participants have more time and possibly can open-source datasets for their own research, and the leaderboard of research gives credit to their methods. For the platform of benchmarks, it is always better to have high-quality benchmarks and attract more people.
@@ 5 Post challenge benchmark
Table 1: Comparisons of competition and benchmark (Xu et al., 2022).
Competition Benchmark
Purpose Crowdsourcing problems in a short time and harvesting solutions Continuous fair evaluation over a long period, in a unified framework
Phases Multiple phases Single-phase
Time period Usually limited Often never-ending
Cooperation & information sharing Limited due to the competitive nature As extensive as possible
Submissions Usually algorithm predictions or algorithm code Algorithm code or datasets; code or dataset name, description, documentation meta-data and fact-sheets; scoring programs for custom analyses
Outcome leaderboard with usually a single global ranking based on one score from each team (last or best) Table with all the submissions made; sorting with multiple scores possible; multiple analyses, graphs, figures, code sharing
6 Conclusion
In the chapter, we have covered the main actions in favour of the long-lasting impact of a challenge, particularly with a post-challenge paper template and advice on how to turn the challenge into a benchmark.
@@ 6 Conclusion
inline,color=brown!40]I commend the authors on a good job at sketching the need for specific activities (such as a paper and benchmarks) for the competition to maximise its impact. Overall, the chapter is in good shape but can benefit from extra work.
@@
Finally, in my view, the chapter reads more like a recipe of how to (best represented by the itemised lists of ingredients) than a piece of work in which the authors share the accumulated wisdom in the field of how to maximise impact with post-challenge activities. In other words, I was hoping for more depth to accompany the breadth. I feel it would be useful, though I leave it to the discretion of the authors, to discuss examples of what different competitions do similarly and differently in terms of post-challenge activities, and how those differences shed light on what works and what doesn’t. In the paper template section, it would be useful to cite references with examples where papers report things in slightly divergent ways (which is good), and papers with sections that are particularly outstanding (e.g., scientific outcomes, ensembling, etc.) in the post-challenge papers in different domains. I know that this may be a tall order, but I honestly believe it will improve the chapter considerably. – Gustavo